# Natal, Brazil

### v1

@@ 5 Discussion
The main goal of our work was to measure developers’ perception regarding the monitoring CI practices in their open-source projects. Our data shows that most developers only consider the automatic execution of the CI build and the automatic execution of tests to define the CI maturity level of the project. They found the seven CI practices presented in our survey important, and they showed interest in monitoring and improving their values. However, as CI services and third-party tools demonstrated only preliminary support to monitor some basic practices, like build status and build duration, there is a lack of a tool that encompasses all CI practices. Furthermore, the use of CI services together with third-party tools may generate unnecessary duplicate monitoring of some CI practices.
We highlight that there is a lack of standardization in the definition and naming of CI (when can also be extended to DevOps) practices, which makes monitoring CI practices more complex. In results of RQ3, we provide a catalog of 32 practices grouped by: (i) CI tools (Figure 14 and Figure 15), (ii) purpose (Figure 17) and (iii) DevOps area (Figure 18). This catalog can be used by future efforts to standardize these practices.
We also conducted an assessment of CI services and third-party tools concerning CI monitoring practices, shown in Figure 16. The data generated by our study can aid in deciding which CI service or third-party tools adopt to monitor these practices as well as highlights which aspects related to monitoring CI practices can be improved in future versions of these tools.
@@
Implications for researchers: Our study showed, especially in RQ2, that developers do not monitor CI practices in a systematic manner and do not use clear criteria to define the maturity level of CI in their projects. We believe there is a wide field to better collect and evaluate the impact of these practices. In addition to defining an approach to better measure the level of CI maturity in projects. Furthermore, this study can be extended to include other DevOps practices.
Implications for practitioners: The results of RQ2 indicate that developers consider the seven CI practices shown in the survey important and would like to monitor them. At the same time, CI tools provide only preliminary support for monitoring CI practices. Our results suggest that there is a need for CI services to implement or improve the monitoring of CI practices. Similarly, standardization of these practices is necessary. These two actions would facilitate and encourage the monitoring of CI practices by project teams. In this sense, initiatives like DORA are very helpful to standardize the definitions.
@@ 6 Threats to Validity
Construct Validity Threats: The lack of standardization in the definition of CI practices across different CI tools can lead to errors in interpreting these practices. In RQ2, survey respondents may respond in a socially desirable way or according to what they think the researcher wants to hear, rather than reflecting their true opinions or feelings.
@@
Although we followed several recommendations ParticipationSurveys2013 to increase the response rate, we received 28 survey responses, resulting in a response rate of 5.6%. However, we consider these responses valuable as they provide insights into the opinions and practices of core developers with extensive knowledge of the CI practice in the analyzed projects. The response rate achieved can be attributed to our focus on selecting qualified developers with expertise in the CI practice. To mitigate the impact of the response rate, we conducted two additional analyses and triangulated the results with those from the survey. To generalize the results for projects of different natures, further replication studies with a larger number of developers are needed.
@@ 7 Conclusion
In this paper, we presented a qualitative study on developers’ perceptions of monitoring CI practices. To achieve this, we conducted a Document Analysis to discover the current needs for monitoring CI practices expressed by developers during the development process, using pull request comments. We then collected the evolution of seven CI practices over a period of 2 years from 121 open-source projects and conducted a survey with 28 developers from these projects. The goal was to understand the importance of the collected practices to the developers and to gain insights into how they measure the adoption of CI in their projects and their needs related to CI monitoring. Finally, we performed another Document Analysis to assess the current level of support provided by CI services for monitoring CI practices.
Our main results were as follows: 1) Coverage is the most monitored DevOps practice, and CodeCov is the most used tool by developers during the development process. 2) There is no standard set of practices or data that developers use to define the maturity level of their CI projects. 3) Most developers found the seven CI practices presented in our survey important and expressed interest in monitoring them if provided the opportunity. 4) CI tools have preliminary support to monitor CI practices, and GitLab is one platform that started to monitor some practices, however with focus on DevOps practices like DORA. 5) We also found a need to monitor “Build health” and “Time to fix a broken build” practices.
Data Availability Statements
The authors declare that they store all data searched and collected in an external public repository available at https://doi.org/10.5281/zenodo.10936371.
Conflict of interest
The authors declare that they have no conflict of interest.

### v2

@@ 5 Discussion
The main goal of our work was to measure developers’ perception regarding the monitoring CI practices in their open-source projects. To achieve this, we calculate CI metrics associated with these practices and look for evidence of monitoring these metrics in PR comments, as well as in the documentation of CI services and third-party tools. Additionally, we conducted a survey with developers to explore the evolution of seven CI metrics in their projects. Our data shows that most developers only consider the automatic execution of the CI build and the automatic execution of tests to define the CI maturity level of the project. They found the seven CI metrics presented in our survey important, and they showed interest in monitoring and improving their values. However, as CI services and third-party tools demonstrated only preliminary support to monitor some basic metrics, like build status and build duration, there is a lack of a tool that encompasses all CI metrics. Furthermore, the use of CI services together with third-party tools may generate unnecessary duplicate monitoring of some CI metrics.
We highlight that there is a lack of standardization in the definition and naming of CI (which can also be extended to DevOps) metrics, which makes monitoring CI metrics more complex. In results of RQ3, we provide a catalog of 32 metrics grouped by: (i) CI tools (Figure 13 and Figure 14), (ii) purpose (Figure 15) and (iii) DevOps area (Figure 16). This catalog can be used by future efforts to standardize these metrics.
We also conducted an assessment of CI services and third-party tools concerning monitoring CI metrics. The data generated by our study can aid in deciding which CI service or third-party tools adopt to monitor these metrics as well as highlights which aspects related to monitoring CI metrics can be improved in future versions of these tools.
Contextual factors such as project size, team structure, and code base complexity undoubtedly play a significant role in influencing the adoption and implementation of Continuous Integration (CI) practices. However, our study deliberately focused on understanding developers’ perceptions of CI monitoring practices, regardless of these contextual variables. This approach allowed us to address a critical yet unexplored gap: the disparity between CI adoption and its proper monitoring. As we have shown, previous studies have highlighted that, even when CI is widely adopted, many projects fail to adequately monitor its use, potentially undermining its intended benefits.
Our work makes unique contributions to the literature on CI, complementing and extending previous studies. First, it offers a broader perspective on CI metrics monitoring and its role in ensuring adherence to CI practices. Using two distinct methods across three research questions to triangulate and reinforce our findings, we gather insights from the perspective of project developers and on a larger number of projects than in previous studies. Specifically, our study highlights two key points: (i) The lack of tools designed for the systematic monitoring and reporting of CI metrics. (ii) Identifying developers’ interest in monitoring CI practices.
The results of our work highlight a clear lack of tools dedicated to monitoring CI metric to ensure that all CI practices are consistently followed by development teams. Our findings also suggest that developers are interested in monitoring these metrics. Consequently, tools that track CI metrics and provide continuous reports to developers could help identify problems and motivate teams to improve their CI process, ultimately improving project quality. Furthermore, our approach can be applied to closed-source or enterprise projects, as long as a monitoring tool can be developed to collect CI metrics in private repositories and adopted by organizations. The monitoring tool can be flexible and configurable so that the team can choose which CI metrics would be collected and the ideal values of these metrics for a specific project. This is the next step in our research.
@@
Implications for researchers: Our study showed, especially in RQ2, that developers do not monitor CI practices in a systematic manner and do not use clear criteria to define the maturity level of CI in their projects. We believe there is a wide field to better collect and evaluate the impact of these practices. In addition to defining an approach to better measure the level of CI maturity in projects or analyzing other aspects of CI practice monitoring not covered by our study, future research could explore how contextual factors, such as project size, team size, and software complexity, influence the monitoring of CI practices. Furthermore, this study can be extended to include other DevOps practices.
Implications for practitioners: Providing clearer strategies, including assigning CI engineers or dedicated monitoring DevOps teams, can help streamline the process and ensure monitoring remains a core part of the development workflow. DevOps teams should incorporate systematic CI monitoring practices into their workflows to ensure continuous feedback on the health of their CI processes. The results of RQ2 indicate that developers consider the seven CI metrics shown in the survey important and would like to monitor them. At the same time, CI tools provide only preliminary support for monitoring CI metrics. Our results suggest that there is a need for CI services to implement or improve the monitoring of CI practices, including increased support for CI metrics monitoring. Effective monitoring tools should be capable of generating automated alerts that notify the team when CI metrics fall below acceptable thresholds, encouraging proactive maintenance of CI health. Regular reports on CI health can further motivate developers by providing a clear view of the project’s state and areas for improvement. Additionally, gamification strategies, such as Competition, Rewards, Feedback Loops for maintaining CI health, can further stimulate developer engagement and encourage consistent best practices. Similarly, standardization of these metrics is necessary. In this sense, initiatives like DORA are very helpful to standardize the definitions. These actions would facilitate and encourage the monitoring of CI practices by project teams.
@@ 6 Threats to Validity
Construct Validity Threats: The lack of standardization in the definition of CI metrics across different CI tools can lead to errors in interpreting these metrics. In RQ2, survey respondents may respond in a socially desirable way or according to what they think the researcher wants to hear, rather than reflecting their true opinions or feelings. To mitigate this threat, we took the following measures: (i) At the beginning of the survey, we explicitly informed participants that all responses would remain anonymous and would be used solely for academic research purposes, (ii) we carefully formulated survey questions to be neutral, avoiding leading language or phrasing that might suggest a “preferred” answer and (iii) we emphasized that participation in the survey was entirely voluntary, and respondents could skip questions or withdraw at any time without any consequences.
@@
A potential limitation of our study is the insufficiency of data for analyzing how contextual factors influence the monitoring of CI practices. While this was not the primary focus of our study, the lack of this data limits the depth of our analysis regarding the broader contextual factors affecting CI monitoring. Future research may aim to collect project-level contextual data to provide a more comprehensive understanding of how these factors influence CI monitoring practices.
Future research should also explore the roles and motivations behind the need for CI practices monitoring within project teams. Understanding who is responsible for tracking CI adoption, enforcing best practices, and interpreting CI metrics can provide valuable insights into how organizations maintain and optimize their CI workflows. This distinction was not explicitly addressed in our study.
@@
Although we followed several recommendations ParticipationSurveys2013 to increase the response rate, we received 28 survey responses, resulting in a response rate of 5.6%. However, we consider these responses valuable as they provide insights into the opinions and practices of core developers with extensive knowledge of the CI practice in the analyzed projects. The response rate achieved can be attributed to our focus on selecting qualified developers with expertise in the CI field. To mitigate the impact of the response rate, we conducted two additional analyses and triangulated the results with those from the survey. To generalize the results for projects of different natures, further replication studies with a larger number of developers are needed.
@@ 7 Conclusion
In this paper, we presented a qualitative study on developers’ perceptions of monitoring CI practices. To achieve this, we conducted a Document Analysis to discover the current needs for monitoring CI practices expressed by developers during the development process, using pull request comments. We then collected the evolution of seven CI metrics associated with CI practices over a period of 2 years from 121 open-source projects and conducted a survey with 28 developers from these projects. The goal was to understand the importance of the collected practices to the developers and to gain insights into how they measure the adoption of CI in their projects and their needs related to CI monitoring. Finally, we performed another Document Analysis to assess the current level of support provided by CI services for monitoring CI practices.
Our main results were as follows: 1) Coverage is the most monitored DevOps metrics, and CodeCov is the most used tool by developers during the development process. 2) There is no standard set of practices or data that developers use to define the maturity level of their CI projects. 3) Most developers found the seven CI metrics presented in our survey important and expressed interest in monitoring them if provided the opportunity. 4) CI tools have preliminary support to monitor CI metrics, and GitLab is one platform that started to monitor some metrics, however with focus on DevOps metrics like DORA. 5) We also found a need to monitor “Build health” and “Time to fix a broken build” metrics.
Declarations
Ethical Approval
Not Applicable
Informed Consent
Informed consent was obtained from all participants prior to completing the survey.
Author Contributions
The first author designed the study, collected data, sent the survey to the participants, and performed the analysis of the results. The second author reviewed the document analysis for RQ1 and RQ3 as well as the manuscript. The third author reviewed the manuscript. The fourth author re-evaluated disagreements between the first and second authors regarding the document analysis of RQ3 and also reviewed the manuscript.
Data Availability Statement
The authors declare that they store all data searched and collected in an external public repository available at https://doi.org/10.5281/zenodo.14569025.
Competing Interests
The authors declare that they have no conflicts of interest or competing interests.
Funding
This work received no funding.